OpenStudy (anonymous):

Two U.S. states are not part of the continental United States. What percent of the fifty U.S. states are included in the continental United States? A. 2% B. 48% C. 96% D. 13%

OpenStudy (ibbutibbu.):

C- 96%

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (ibbutibbu.):

well, \[\frac{50}{50} - \frac{ 2 }{ 50 } = ?\]

OpenStudy (ibbutibbu.):

\[\frac{ 48 }{ 50 } = \frac{ x }{ 100 }\]

OpenStudy (surana):

96%

OpenStudy (ibbutibbu.):

@Surana you cant give answers

OpenStudy (ibbutibbu.):

you need to explain

OpenStudy (surana):

Oh, I'm sorry, I didn't know that.

OpenStudy (ibbutibbu.):

no problem :)

OpenStudy (surana):

Okay, the reasoning behind why I chose 96% is because 50 states could be constituted as 100%, 1 state constitutes 2 percent, and 25 states constitute 50%. 2 states therefore constitute 4%, and subtracting 4 from 100 leaves us with 96, hence the choice for 96%, which is C.

OpenStudy (ibbutibbu.):

Better :)

OpenStudy (surana):

Glad to know! As this is my second day here I had no idea what to do and what not to do, but oh well. Glad to have gotten some help.

OpenStudy (ibbutibbu.):

Need anything, just tell me!

OpenStudy (surana):

Thanks! I do have a couple of problems regarding math to ask, so I'll head over and work with those.

OpenStudy (anonymous):

So if two states are not included in the continental United States, and they want to know the percentage of states that are included in the continental United States, then: \(\sf\Large\frac{Number~of~States~in~Continental~United~States}{Total~States~in~United~States}\) And I'm pretty sure you can figure it out from there. :)